Solve the following equation. Solve for X.
-x/10 = =-8/3

Answer by stanbon(75887)

Solve for X.
-x/10 = =-8/3
--------
x/10 = 8/3
x = 10(8/3)
---
x = 26 2/3
==========================
Cheers,
Stan H.
==========================